# Structural dynamics and optimal transport of an active polymer

### v3

@@ Abstract
where $f(\psi)=-d\Phi(\psi)/d\psi$ is an effective “force”, $x=\psi/U$ and $\tau=t/\mathcal{T}$ with $U^{2}=2a_{4}/(3a_{6})$, $\mathcal{T}=3a_{6}/(8a^{2}_{4})$ and $\chi=-3a_{2}a_{6}/(4a^{2}_{4})$. The values of $a_{2}$, $a_{4}$ and $a_{6}$ as a function of $F_{\textrm{sp}}$ and the corresponding values of $\chi$, $\mathcal{T}$ and $U$ can be found in Table I in the SM. The term $\xi$ refers to a white noise that satisfies $\langle\xi(\tau)\rangle=0$ and $\langle\xi(\tau)\xi(\tau^{{}^{\prime}})\rangle=2\Gamma\delta(\tau-\tau^{{}^{\prime}})$. The noise intensity $\Gamma$ is obtained by mapping the PDF of the turning number in LD simulations with the one resulting from the integration of Eq. (2) [see Fig. 3(a)] and has a fixed value of $\Gamma=0.00375$ for different values of $F_{\textrm{sp}}$. The deterministic form of Eq. (2) – that corresponds to $\Gamma=0$ and thus $\dot{x}=f(x;\chi)$ – undergoes a subcritical pitchfork bifurcation strogatz . Time evolution of the solution of Eq. (2) for different values of the initial normalized turning number $x_{0}$ and for fixed value of $\chi=-0.184$ corresponding to $F_{\textrm{sp}}=1.9$ can be seen in the SM. The bifurcation diagram in the $x-\chi$ plane is shown in Fig. 3(b). The fixed points of Eq. (2) – $f(x^{*};\chi)=0$ – are:
@@
For $\chi>0$, $x^{*}=0$ is an (linearly) unstable fixed point since $f^{{}^{\prime}}(x^{*})=\frac{\partial f}{\partial x}(x^{*})\geq 0$. For $\chi\leq 0$, $x^{*}=0$ becomes stable ($f^{{}^{\prime}}(0)<0$) and two unstable fixed points – such that $f^{{}^{\prime}}(x^{*})\geq 0$ – emerge. At $\chi_{\textrm{s}}=-1/4$ there are two additional saddle-node bifurcations [denoted by a black star in the main panel of Fig. 3(b)] which are obtained by considering the condition $1+4\chi_{\textrm{s}}\geq 0$. Note that the fixed points are the extremum values of the $P(\psi)$ [see Fig. 1(b) in SM]. The effect of $x^{5}$ term in Eq. (2) in the bifurcation diagram is to turn around the unstable branches at $\chi=\chi_{\textrm{s}}$ and become stable for $\chi>\chi_{\textrm{s}}$ [solid blue lines in the top and bottom in Fig. 3(b) main panel]. These stable states provide the possibility of jumps and hysteresis as $\chi$ is varied, that is the characteristic of the first-order phase transitions. The filled and empty green circles in Fig. 3(b) correspond to the LD simulations and show the stable and unstable states, respectively. The simulations data and theoretical curves show a perfect agreement. The inset in Fig. 3(b) shows how the values of $F_{\textrm{sp}}$ and $\chi$ relate to each other.
@@
To compute the diffusion coefficient $D_{\textrm{CM}}$, and to put in evidence the run-&-tumble character of the process, we consider that during the spiral states, the sum of all active forces acting on different segments of the polymer is negligible. Thus, in this state, the speed (of the CM) of polymer is $v_{\textrm{s}}\simeq 0$. This implies that we assume that the distance travelled by the polymer in the spiral sates is negligible [see the movie in SM]. On the other hand, in the non-spiral states, the polymer is (in comparison to spiral states) extended, and thus, the sum of active forces is non-zero. In consequence, in the non-spiral state, the polymer moves at speed $v_{\textrm{ns}}>0$, and we assume that it travels an average distance $\ell=v_{\textrm{ns}}\langle\tau_{\textrm{ns}}\rangle$. To determine $v_{\textrm{ns}}$, let us consider that in the non-spiral state, the polymer is an extended rod – that experiences an active force $NF_{\textrm{sp}}/\ell_{\textrm{p}}$ and a friction drag $N\,\gamma$ – and thus, moves at speed $v_{\textrm{ns}}=F_{\textrm{sp}}/\gamma~{}\!\ell_{\textrm{p}}$. All together, this implies that as the polymer goes through a cycle spiral-non-spiral, i.e. involving a transition from the spiral to the non-spiral state, moves an average distance $\ell$. Since the average duration of these cycles is $\Omega=\langle\tau_{\textrm{s}}\rangle+\langle\tau_{\textrm{ns}}\rangle$, in a time $t$, the polymer performs an average of $n=t/\Omega$ of those cycles. It is worth noting that $\langle\tau_{\textrm{ns}}\rangle$ and $\langle\tau_{\textrm{s}}\rangle$ can be estimated from the transition rates to escape from the non-spiral and spiral states considering the associated Kramer’s escape problem that assumes these rates are proportional to the exponential of the potential barrier: $\langle\tau_{\textrm{ns}}\rangle^{-1}\propto\exp(\Delta\Phi_{1}/\Gamma)$ and $\langle\tau_{\textrm{s}}\rangle^{-1}\propto\exp(\Delta\Phi_{2}/\Gamma)$, respectively, where $\Delta\Phi_{1}=\frac{1}{12}(1+4\chi)^{3/2}$ and $\Delta\Phi_{2}=\frac{1}{48}(-1+\sqrt{1+4\chi})(1+8\chi-\sqrt{1+4\chi})$ [see Fig. 3(c) bottom panel]. The detail of how $\Delta\Phi_{1}$ and $\Delta\Phi_{2}$ are obtained, can be seen in the SM. Recall that then $p_{\textrm{s}}$ and $p_{\textrm{ns}}$ can also be expressed in term of these rates, and that $\chi=\chi(F_{\textrm{sp}})$ [see Fig. 3(d)]. Finally, let us recall that for a random walk (RW) after $n_{\textrm{RW}}$ steps of length $\ell_{\textrm{RW}}$, the mean-square displacement goes as $\langle\mathbf{r}^{2}\rangle=n_{\textrm{RW}}\,\ell_{\textrm{RW}}^{2}$. Thus, if we interpret that in time $t$, the $n$ cycles performed by the active polymer correspond to $n_{\textrm{RW}}$ steps of a RW, that moves in each step an average distance $\ell_{\textrm{RW}}=\ell$, we can approximate the diffusion coefficient – using the above-given expressions for $n$ and $\ell$ – as:
@@
where $\overline{v}=p_{\textrm{ns}}v_{\textrm{ns}}$, and is in agreement with the run-and-tumble particle model Sokolov_PRE_2012 . The comparison between the diffusion coefficients $D_{\textrm{CM}}$ predicted by Eq. (4) (denoted by ”Theory”; orange squares) and the one obtained from the LD simulations (”LD Sim”; black circles) displayed in Fig. 2(a), shows that there exists a good quantitative agreement.
@@
Importantly, Eqs. (2) and (4) allow us to prove that there is an optimal self-propelling force that maximizes the diffusion coefficient. The existence of an optimal self-propelling force is easy to understand. When $F_{\textrm{sp}}\to 0$, the spiral states is not observed. The speed in the non-spiral state is $v_{0}\propto F_{\textrm{sp}}$, then, $F_{\textrm{sp}}\to 0$ implies $D_{\textrm{CM}}\to 0$. On the contrary, for $F_{\textrm{sp}}\gg 1$, the polymer is locked in spirals, but since we assume that $v_{0}=0$ in this state, we get again  $D_{\textrm{CM}}\to 0$. In summary, increasing $F_{\textrm{sp}}$ increases $v_{0}$, but at the same time enhances the probability of finding the polymer in the spiral state, where the motion of the center of mass of the polymer is strongly reduced. And thus, we find that for intermediate values of $F_{\textrm{sp}}$, $D_{\textrm{CM}}$ is maximized. Compare the results obtained here with those reported in a numerical study with the Krotky-Porod model theo4 .
@@
Our studies puts in evidence the existence of a key coupling between the structural dynamics of the polymer and its transport properties. Moreover, we have shown that there exist an optimal active force $F_{\textrm{sp}}$ that maximizes the diffusion coefficient $D_{\textrm{CM}}$. The reported results are of relevance for a large number of experimental active polymer studies, including actin filament motion in motility assays experiments and DNA duplication, among many other examples.
Acknowledgements.
FP acknowledges fruitful discussions with Debasish Chaudhury.

### v4

@@ IV Theory
where $f(x)=-d\Phi(x)/dx$ is an effective “force”, $x=\psi/U$ and $\tau=t/\mathcal{T}$ with $U^{2}=2a_{4}/(3a_{6})$, $\mathcal{T}=3a_{6}/(8a^{2}_{4})$ and $\chi=-3a_{2}a_{6}/(4a^{2}_{4})$. The values of $a_{2}$, $a_{4}$ and $a_{6}$ as a function of $F_{\textrm{sp}}$ and the corresponding values of $\chi$, $\mathcal{T}$ and $U$ can be found in Table 1.
Table 1: Fitting parameters for $\ell_{p}=5$.
$F_{\textrm{sp}}$ $a_{2}$ $a_{4}$ $a_{6}$ $\chi$ $\mathcal{T}$ $U$
0.0 1.337 0.00348 0.00161 -132.947 49.715 1.200
0.5 0.372 0.0205 0.000534 -0.353 0.473 5.065
1.0 0.307 0.0164 0.0002729 -0.231 0.376 6.344
1.25 0.362 0.0213 0.000342 -0.204 0.282 6.443
1.5 0.447 0.0272 0.000424 -0.192 0.214 6.539
1.75 0.569 0.0346 0.000523 -0.186 0.163 6.641
1.9 0.736 0.0444 0.000658 -0.184 0.125 6.706
2.0 0.857 0.0512 0.000749 -0.183 0.107 6.751
2.25 1.153 0.0677 0.000966 -0.182 0.0789 6.837
2.5 1.233 0.0716 $0$0.180 0.0733 6.900
3.0 0.823 0.0492 0.000684 -0.174 0.105 6.927
4.0 0.497 0.0336 0.000476 -0.1573 0.158 6.856
5.0 0.490 0.0351 0.000510 -0.151 0.154 6.776
8.0 0.45 0.0350 0.00050 -0.137755 0.153 6.831
10.0 -0.150 0.0350 0.00050 0.0459 0.1531 6.831
13.0 -0.165 0.0350 0.00050 0.05051 0.15306 6.831
15.0 -0.180 0.0350 0.00050 0.0551 0.1531 6.831
20.0 -0.195 0.0350 0.00050 0.0597 0.1531 6.831
The term $\xi$ refers to a white noise that satisfies $\langle\xi(\tau)\rangle=0$ and $\langle\xi(\tau)\xi(\tau^{{}^{\prime}})\rangle=2\Gamma\delta(\tau-\tau^{{}^{\prime}})$. The noise intensity $\Gamma$ is obtained by mapping the PDF of the turning number in LD simulations with the one resulting from the integration of Eq. (3) [see Fig. 4(a)] and has a fixed value of $\Gamma=0.00375$ for different values of $F_{\textrm{sp}}$. Note that though the noise in Eq. (3) was assumed white, a colored noise is, arguably, a more generic option. In short, the dynamics of $\psi$ can be expressed, instead of using Eq. (3), by a generalized Langevin equation with the memory kernelPanja_J_Stat_Mech ; Vandebroek_J_Stat_Phys and/or a mobility kernel Sakaue_PRE_2013 ; Saito_PRE_2015 , or a combination of both Shinkai . Let us recall that memory effects occur, for instance, in protein folding Polymer_folding_PRE_2102 ; Polymer_folding_PRL_2104 ; Soft_Matter_2017 and in polymer translocation Poly_trans_J_Cond_Matt_2013 ; Poly_trans_Soft_Matter_2014 ; Poly_trans_Polymers_2016 , where tension propagation theory provides a reliable description Sakaue_PRE_2007 ; Rowghanian_2011 ; Timo_PRE_2012 ; Jalal_JCP_2014 ; Jalal_JCP_2015 ; Jalal_EPL_2017 ; Jalal_Sci_Rep_2017 ; Jalal_Cond_Matt_2018 ; Jalal_Cond_Matt_2020 ; Jalal_PRR_2021 ; Jalal_PRR_2023 . However, for the dynamics of $\psi$, as we show below, it is sufficient with Eq. (3) and a white noise: Eq. (3) does not only allow us to obtain the correct steady state of $P(\psi)$, it also captures the correct transition rates between spiral and non spiral configurations.
The deterministic form of Eq. (3) – that corresponds to $\Gamma=0$ and thus $\dot{x}=f(x;\chi)$ – undergoes a subcritical pitchfork bifurcation strogatz ; the time-dependent solution of Eq. (3) is shown in Appendix D. The bifurcation diagram in the $x^{*}-\chi$ plane is shown in Fig. 4(b). The fixed points of Eq. (3) – $f(x^{*};\chi)=0$ – are:
@@
For $\chi>0$, $x^{*}=0$ is an (linearly) unstable fixed point since $f^{{}^{\prime}}(x^{*})=\frac{\partial f}{\partial x}(x^{*})\geq 0$. For $\chi\leq 0$, $x^{*}=0$ becomes stable ($f^{{}^{\prime}}(0)<0$) and two unstable fixed points – such that $f^{{}^{\prime}}(x^{*})\geq 0$ – emerge. At $\chi_{\textrm{s}}=-1/4$ there are two additional saddle-node bifurcations [denoted by violet stars in the main panel of Fig. 4(b)] which are obtained by considering the condition $1+4\chi_{\textrm{s}}\geq 0$. Note that the fixed points are the extremum values of the $P(\psi)$ [see Fig. 6 in the Appendix B]. The effect of $x^{5}$ term in Eq. (3) in the bifurcation diagram is to turn around the unstable branches at $\chi=\chi_{\textrm{s}}$ and become stable for $\chi>\chi_{\textrm{s}}$ [solid blue lines in the top and bottom in Fig. 4(b) main panel]. These stable states provide the possibility of jumps and hysteresis as $\chi$ is varied. The filled and empty green circles in Fig. 4(b) correspond to the LD simulations and show the stable and unstable states, respectively. The simulations data and theoretical curves show a perfect agreement. The inset in Fig. 4(b) shows how the values of $F_{\textrm{sp}}$ and $\chi$ relate to each other.
@@
To compute the diffusion coefficient $D_{\textrm{CM}}$, and to put in evidence the run-&-tumble character of the process, we consider that during the spiral states, the sum of all active forces acting on different segments of the polymer is negligible. Thus, in this state, the speed (of the CM) of polymer is approximately $v_{\textrm{s}}\simeq 0$. This implies that we assume that the distance travelled by the polymer in the spiral sates is negligible [see the movie in SM]. On the other hand, in the non-spiral states, the polymer is (in comparison to spiral states) extended, and thus, the sum of active forces is non-zero. In consequence, in the non-spiral state, the polymer moves at speed $v_{\textrm{ns}}>0$, and we assume that it travels an average distance $\ell=v_{\textrm{ns}}\langle\tau_{\textrm{ns}}\rangle$. To approximately determine $v_{\textrm{ns}}$, let us assume that in the non-spiral state, the polymer is an extended rod – that experiences an active force $NF_{\textrm{sp}}/\ell_{\textrm{p}}$ and a friction drag $N\,\gamma$ – and thus, moves at speed $v_{\textrm{ns}}=F_{\textrm{sp}}/\gamma~{}\!\ell_{\textrm{p}}$. All together, this implies that as the polymer goes through a cycle spiral-non-spiral, i.e. involving a transition from the spiral to the non-spiral state, moves an average distance $\ell$. Since the average duration of these cycles is $\Omega=\langle\tau_{\textrm{s}}\rangle+\langle\tau_{\textrm{ns}}\rangle$, in a time $t$, the polymer performs an average of $n=t/\Omega$ of those cycles. It is worth noting that $\langle\tau_{\textrm{ns}}\rangle$ and $\langle\tau_{\textrm{s}}\rangle$ can be estimated from the transition rates to escape from the non-spiral and spiral states considering the associated Kramer’s escape problem that assumes these rates are proportional to the exponential of the potential barrier: $\langle\tau_{\textrm{ns}}\rangle^{-1}=A_{\textrm{ns}}\exp(\Delta\Phi_{1}/\Gamma)$ and $\langle\tau_{\textrm{s}}\rangle^{-1}=A_{\textrm{s}}\exp(\Delta\Phi_{2}/\Gamma)$, respectively van_Kampen , where $\Delta\Phi_{1}=\frac{1}{12}(1+4\chi)^{3/2}$ and $\Delta\Phi_{2}=\frac{1}{48}(-1+\sqrt{1+4\chi})(1+8\chi-\sqrt{1+4\chi})$ [see Fig. 4(c) bottom panel]. The detail of how $\Delta\Phi_{1}$ and $\Delta\Phi_{2}$ are obtained, can be seen in the Appendix D. The estimates for $A_{\textrm{ns}}$ and $A_{\textrm{ns}}$ in Fig. 3(c) are $A_{\textrm{ns}}=A_{\textrm{s}}=2.35$. Recall that then $p_{\textrm{s}}$ and $p_{\textrm{ns}}$ can also be expressed in term of these rates: $p_{\textrm{s}}=\langle\tau_{\textrm{s}}\rangle/\Omega$ and $p_{\textrm{ns}}=\langle\tau_{\textrm{ns}}\rangle/\Omega$. This is shown in Fig. 4(d). Note that there is a good agreement between estimates from LD simulations and these theoretical estimates.
Finally, let us recall that for a random walk (RW) after $n_{\textrm{RW}}$ steps of length $\ell_{\textrm{RW}}$, the MSD is written as $\langle\mathbf{r}^{2}\rangle=n_{\textrm{RW}}\,\ell_{\textrm{RW}}^{2}$ Reif . Thus, if we interpret that in time $t$, the $n$ cycles performed by the active polymer correspond to $n_{\textrm{RW}}$ steps of a RW, that moves in each step an average distance $\ell_{\textrm{RW}}=\ell$, we can approximate the diffusion coefficient – using the above-given expressions for $n$ and $\ell$ – as:
@@
where $\overline{v}=p_{\textrm{ns}}v_{\textrm{ns}}$, and is in agreement with the run-and-tumble particle model Sokolov_PRE_2012 . The comparison between the diffusion coefficients $D_{\textrm{CM}}$ predicted by Eq. (5) (denoted by ”Theory”; orange squares) and the one obtained from the LD simulations (”LD Sim”; black circles) displayed in Fig. 3(c), shows that there exists a good qualitative agreement between LD simulations and the theory.
@@
Importantly, Eqs. (3) and (5) allow us to prove that there is an optimal self-propelling force that maximizes the diffusion coefficient. The existence of such an optimal force value is easy to understand. When $F_{\textrm{sp}}\to 0$, the spiral states is not observed. The speed in the non-spiral state is $v_{0}\propto F_{\textrm{sp}}$, then, $F_{\textrm{sp}}\to 0$ implies $D_{\textrm{CM}}\to 0$. On the contrary, for $F_{\textrm{sp}}\gg 1$, the polymer is locked in spirals, but since we assume that $v_{0}~{}\!\simeq~{}0$ in this state, we get again  $D_{\textrm{CM}}\to 0$. In summary, increasing $F_{\textrm{sp}}$ increases $v_{0}$, but at the same time enhances the probability of finding the polymer in the spiral state, where the motion of the center of mass of the polymer is strongly reduced. And thus, we find that for intermediate values of $F_{\textrm{sp}}$, $D_{\textrm{CM}}$ is maximized. Compare the results obtained here with those reported in a numerical study with the Krotky-Porod model theo4 .
V Conclusions
@@
Our study puts in evidence the existence of a key coupling between the structural dynamics of the polymer and its transport properties. Moreover, we have shown that there exists an optimal active force $F_{\textrm{sp}}$ that maximizes the diffusion coefficient $D_{\textrm{CM}}$. It would be interesting to employ the optimal Langevin modeling of out-of-equilibrium molecular dynamic simulations and compare it with the approach developed here OptimalLangevin . The reported results are of relevance for a large number of experimental active polymer studies, including actin filament motion in motility assays experiments, among many other examples.
Figure 5: (a) The absolute value of the maximum turning number $|\psi_{m}|$ as a function of $F_{\textrm{sp}}$ (blue circles) corresponding to the left and right peaks in Fig. 6. Inset shows the value of $B$ obtained from the Eq. (8) as a function of $F_{\textrm{sp}}$ by employing the values of the $|\psi_{m}|$ from the main panel. (b) The radius of gyration of the polymer $R_{\textrm{g}}$ as a function of $F_{\textrm{sp}}$ for fixed value of the persistence length $\ell_{p}=5$, and for mixture of both non-spiral and spiral states (denoted by all states, blue circles), for just spiral state (green squares), and obtained from Eq. (7) by the theory (red dashed line). Inset shows the Archimedean spiral.
VI Appendix A: Radius of gyration and turning number; theory & simulation
The spatial size of the polymer in the spiral state, specifically radius of gyration (RG), is related to $\psi$. Assume that the polymer can be approximately described by an Archimedean spiral [inset of Fig. 5(b)] in which the radius $r$ is a function of the angle $\varphi$ such that $r(\varphi)=A+B\varphi/(2\pi)$. The coefficients $A$ and $B$ are functions of $F_{\textrm{sp}}$ and $\ell_{\textrm{{p}}}$. Under this assumption, the RG of polymer $R_{\textrm{g}}$ with contour length of $N$ is computed as  rubin_stein
[…]
$R_{\textrm{g}}=\bigg{\{}\frac{\pi}{2BN}\big{[}\big{(}A+B\psi_{\textrm{m}}\big{)}^{4}-A^{4}\big{]}\bigg{\}}^{1/2}\,,$ (7)
[…]
$\pi B\psi_{\textrm{m}}^{2}+2\pi A\psi_{\textrm{m}}-N=0.$ (8)
Then, $\psi_{\textrm{m}}$ is obtained from the roots of Eq. (8). In the main panel (a) in Fig. 5 the absolute value of the maximum turning number $|\psi_{\textrm{m}}|$ has been plotted as a function of $F_{\textrm{sp}}$, that are obtained from the LD simulations by considering the most probable values of $\psi$ (left and right peaks in Fig. 6). The inset shows the value of the coefficient $B$ as a function of $F_{\textrm{sp}}$. In Fig. 5(b) using the LD simulations the $R_{\textrm{g}}$ has been plotted as a function of $F_{\textrm{sp}}$ when contributions of both non-spiral and spiral states have been taken into account (blue circles) and also when only the spiral state has been taken into account (green diamonds). As seen, the RG is a monotonic and decreasing function of the SP force due to the increasing of probability of finding the polymer in the spiral state [see Fig. 3(d)]. Substituting the values of $A=1$ and $B$ [presented in the inset of Fig. 5(a)] into Eq. (7), the semi-analytical values of the RG as a function of SP force is plotted in Fig. 5(b) (red dashed line). As seen the semi-analytical data are in a very good agreement with those coming from the LD simulations for the spiral state.
VII Appendix B: PDF of turning number
Fig. 6 represents the PDF of the turning number $P(\psi)$ for an active polymer with contour length of $N=200$, persistence length of $\ell_{\textrm{p}}=5$, for various values of the SP forces $F_{\textrm{sp}}=$0.0 (blue circles), 1.0 (red squares), 1.9 (green diamonds), 3.0 (pink triangles-up) and 5.0 (orange triangles left).
Figure 6: The PDF of the turning number $P(\psi)$ with fixed values of the polymer contour length of $N=200$, persistence length of $\ell_{\textrm{p}}=5$, for various values of of the SP forces $F_{\textrm{sp}}=$0.0 (blue circles), 1.0 (red squares), 1.9 (green diamonds), 3.0 (pink triangles-up) and 5.0 (orange triangles left).
VIII Appendix C: Transitions between the non-spiral and the spiral states
In general, there are three different types of transitions between the non-spiral and the spiral states. In panel (a) of Fig. 7 the time evolution of the turning number $\psi(t)$ has been plotted. Panels (b) and (c) are devoted to present the corresponding speed of the polymer CM $|v(t)|$ and displacement of the polymer CM $|\Delta(t)|$, respectively. Three different types of the transitions are distinguished by the three vertical color boxes. The type $1$ (cyan box) is a spiral-non-spiral-spiral transition, in which the polymer transients from a CW spiral state to a non-spiral state, and then to a CCW spiral state. The inverse transition is also possible, i.e. from a CCW spiral state to a non-spiral state, and then to a CW spiral state (not shown here). In type $2$ (orange box), the polymer is opened up from a CCW spiral state to a non-spiral state, and then gets back to a CCW spiral state. Obviously, the transition CW-non-spiral-CW is also possible. In type $3$ (magenta box), the polymer directly transients from a CCW spiral state to a CW spiral state without passing from an intermediate non-spiral state, again with possibility of the inverse transition. In type $3$, the polymer starts to be opened up from its interior layers and changes its direction of rotation without any passing through an intermediate non-spiral transient state. In panels (b) and (c) of Fig. 7, the peaks represent the transition to the non-spiral state. As seen, there is not any peak for the type $3$ transition due to the direct transition between CCW and CW spiral states without any passing through an intermediate non-spiral transient state. To show the three types of the transition a typical movie has been provided in the SM.
Figure 7: Three types of transition between the non-spiral and the spiral states. (a) Turning number as a function of time. Panels (b) and (c) are the speed and displacement of the polymer CM, respectively, as a function of time corresponding to panel (a). Each type of transition is distinguished by a different color box (see the text in the Appendix C for more details). The values of the system parameters are $N=200$, $\ell_{\textrm{p}}=5$ and $F_{\textrm{sp}}=4$.
Figure 8: (a) Time evolution of the solution (the normalized turning number $x$) of Eq. (9) for different values of the initial normalized turning number $x_{0}$ and for fixed value of $\chi=-0.184$ corresponding to $F_{\textrm{sp}}=1.9$. The solid and dashed blue lines represent the stable and unstable fixed points corresponding to the $\chi=-0.184$, respectively. (b) The potential $\Phi(x)$ as a function of $x$ for $\chi=-0.2$. The filled and open green circles represent the stable and unstable extrema (fixed points) of the $\Phi(x)$, respectively. The $\Delta\Phi_{1}$ and $\Delta\Phi_{2}$ are the potential barriers that the polymer has to overcome to make transitions between spiral and non-spiral states.
IX Appendix D: Bifurcation and potential barriers
The deterministic form of the time evolution of the normalized turning number $x$ is described by the following relation (see also the text in Sec. IV for more details)
$\frac{dx}{d\tau}=f(x)=\chi x+x^{3}-x^{5}.$ (9)
Fig. 8(a) shows that the normalized turning number $x$ asymptotically approaches to a stable fixed point depending on the initial value of $x$. Fig. 8(a) presents the numerical solutions of Eq. (9) for fixed value of $\chi=-0.184$ corresponding to $F_{\textrm{sp}}=1.9$, and for various initial values of the normalized turning number $x_{0}=2.0$ (in black color), $0.5$ (in red color), $0.2$ (in green color), $-0.2$ (in orange color), $-0.5$ (in brown color) and $-2.0$ (in pink color). The blue solid lines represent the stable fixed points at $x^{*}=0.87,0.0$ and $-0.87$, while the blue dashed lines show the unstable fixed points at $x^{*}=0.493$ and $-0.493$.
Fig. 8(b) shows the potential $\Phi(x)$ as a function of $x$ for fixed value of $\chi=-0.2$. The filled green circles (located at $x^{*}_{1},x^{*}_{3}$ and $x^{*}_{5}$) and open green circles (located at $x^{*}_{2}$ and $x^{*}_{4}$) represent the stable and unstable fixed points of the system, respectively, that are given by solving $f(x^{*})=-\partial\Phi(x)/\partial x|_{x^{*}}=0$, as $x^{*}_{1}=-[(1+\sqrt{1+4\chi})/2]^{1/2},x^{*}_{2}=-[(1-\sqrt{1+4\chi})/2]^{1/2},x^{*}_{3}=0,x^{*}_{4}=[(1-\sqrt{1+4\chi})/2]^{1/2}$ and $x^{*}_{5}=[(1+\sqrt{1+4\chi})/2]^{1/2}$. To make transitions between the spiral and non-spiral states, the polymer has to overcome the potential barriers denoted by the $\Delta\Phi_{1}$ and $\Delta\Phi_{2}$ in Fig. 8(b). It is noticeable that the region between $x^{*}_{2}$ and $x^{*}_{4}$ corresponds to the spiral state and otherwise to the non-spiral state (for more detail see the text in Sec. IV). As it can be seen in Fig. 8(b), the potential barriers are given by $\Delta\Phi_{1}=|\Phi(x^{*}_{2})-\Phi(x^{*}_{1})|=|\Phi(x^{*}_{4})-\Phi(x^{*}_{5})|$ and $\Delta\Phi_{2}=|\Phi(x^{*}_{2})-\Phi(x^{*}_{3})|=|\Phi(x^{*}_{4})-\Phi(x^{*}_{3})|$, where $\Phi(x)=-\chi x^{2}/2-x^{4}/4+x^{6}/6$. Therefore, the potential barriers are written as $\Delta\Phi_{1}=\frac{1}{12}(1+4\chi)^{3/2}$ and $\Delta\Phi_{2}=\frac{1}{48}(-1+\sqrt{1+4\chi})(1+8\chi-\sqrt{1+4\chi})$.
Acknowledgements
FP acknowledges fruitful discussions with Debasish Chaudhury. JS acknowledges Iran National Science Fundation ”This work is based upon research funded by Iran National Science Fundation (INSF) under project No. 4026895”.